OopsSorry1:

which of these expressions is equivalent to log (4^6)? here are the options: a: log(6)*log(4) b: 6*log(4) c: log(6)-log(4) d: log(6)+log(4) Let me know what you all think about this

Coyote:

Well ^6 is an exponent so that should help you out a little.

Coyote:

So you're not adding or subtracting... and definitely not multiplying both of them together.

Coyote:

Which means your answer comes down to B. Hope this helps!

m4rs:

I agree with Coyote, your answer should come down to B.

Coyote:

There should be a button right beside the medal icon that says "Best Response", can you click it?

jhonyy9:

@oopssorry1 just you need know the properties of logarithms log(a) b = c => a^c = b log a + log b = log a×b log a - log b = log a:b log a^x = xlog a so this last one is what you need use in case of your posted question